celticcat:

Evaluating (-0.3) / (-0.2) + (-2.56) = -1.06 is correct right. My daughter says her teacher says the answer is -2.65!! They did the addition first...

Vocaloid:

Yeah -1.06 is the correct solution for the problem as written No idea what her teacher is smoking

Nnesha:

\[\rm \frac{ -0.3 }{ -0.2+(-2.56) } ~~or~~\frac{ -0.3 }{ -0.2 }+(-2.56)\] there is no way to get the answer -2.56..doesn't matter if we add or divide first.

justjm:

teacher deserves a neck on this one

jhonyy9:

,, They did the addition first.." so in this case assume -0,2 to -2,56 and this will be the denominator -2,76 so in this case dividing minus by minus not can getting minus like result just plus (-0,3)/(-0,2) = 1,5 1,5 +(-2,56) = - 1,06